Let L: R^2 --> R be a function defined by L(x,y)=x^2 + y^2 . L(2,3)+L(3,4)=?

a. 38

b. 32

c. 34

d. 36
Answer #1

Hey there stay home and stay safe

\small \\$Given that $\\ L(x,y)=x^2+y^2\\ L(2,3)+L(3,4)\\ =2^2+3^2+3^2+4^2\\ =4+9+9+16\\ =38

So a is the right answer

And please leave a like. Thank you.